# Finite simple groups acting with fixity $4$ and their occurrence as groups of automorphisms of Riemann surfaces

### v1

@@ Abstract.
In previous work, all finite simple groups that act with fixity 4 have been classified. In this article we investigate which ones of these groups act faithfully on a compact Riemann surface of genus at least $2$ with fixity four in total and in such a way that fixity 4 is exhibited on at least one orbit.
@@ 1 Introduction
In the following table, $q$ always denotes a prime power. We also remark that in particular $g_{0}\geq 0$ if a Hurwitz datum has at least three branchings, and otherwise $g_{0}\geq 1$. But there are some exceptions which we mention in the Remark column.
In Lines 8a–8i we set $\alpha:=q^{2}+\sqrt{2}q+1$ and $\beta:=q^{2}-\sqrt{2}q+1$, and $q$ is a power of $2$.
@@
Line Hurwitz datum Remark
1a $[\textrm{Alt}_{7},g,g_{0}\mid[5,1]]$
1b $[\textrm{Alt}_{7},g,g_{0}\mid[5,1],[7,1]]$
2a $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1]]$ $g_{0}\geq 2$
2b $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,1]]$
2c $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,2]]$ $g_{0}\geq 1$
2d $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[7,1]]$
2e $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,1],[7,1]]$
2f $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,2],[7,1]]$
3a $[\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1]]$ $g_{0}\geq 2$
3b $[\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,1]]$
3c $[\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,2]]$
3d $[\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[9,1]]$
3e $[\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[9,2]]$
3f $[\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,1],[9,1]]$
3g $[\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,2],[9,1]]$
3h $[\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,1],[9,2]]$
3i $[\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,2],[9,2]]$
4a $[\textrm{PSU}_{4}(3),g,g_{0}\mid[5,1]]$
4b $[\textrm{PSU}_{4}(3),g,g_{0}\mid[5,1],[7,1]]$
5a $[\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q-1}{4},1]]$ $q\geq 9$, $q\equiv 1\mod 4$
5b $[\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q+1}{4},1]]$ $q\geq 9$, $q\equiv-1\mod 4$
5c $[\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q-1}{4},1],[\frac{q+1}{2},1]]$ $q\geq 9$, $q\equiv 1\mod 4$
5d $[\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q-1}{4},1],[\frac{q+1}{2},2]]$ $q\geq 9$, $q\equiv 1\mod 4$
5e $[\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q+1}{4},1],[\frac{q-1}{2},1]]$ $q\geq 9$, $q\equiv-1\mod 4$
5f $[\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q+1}{4},1],[\frac{q-1}{2},2]]$ $q\geq 9$, $q\equiv-1\mod 4$
6 $[\textrm{PSp}_{4}(q),g,g_{0}\mid[\frac{q^{2}+1}{(2,q^{2}-1)},1]]$ $q\geq 3$
7 $[\textrm{P$\Omega$}^{-}_{8}(q),g,g_{0}\mid[\frac{q^{4}+1}{(2,q^{4}-1)},1]]$
8a $[\textrm{Sz}(q),g,g_{0}\mid[\alpha,1]]$ $q\geq 8$
8b $[\textrm{Sz}(q),g,g_{0}\mid[\beta,1]]$ $q\geq 8$
8c $[\textrm{Sz}(q),g,g_{0}\mid[\alpha,1],[\beta,1]]$ $q\geq 8$
8d $[\textrm{Sz}(q),g,g_{0}\mid[\alpha,1],[q-1,1]]$ $q\geq 8$
8e $[\textrm{Sz}(q),g,g_{0}\mid[\alpha,1],[q-1,2]]$ $q\geq 8$
8f $[\textrm{Sz}(q),g,g_{0}\mid[\beta,1],[q-1,1]]$ $q\geq 8$
8g $[\textrm{Sz}(q),g,g_{0}\mid[\beta,1],[q-1,2]]$ $q\geq 8$
8h $[\textrm{Sz}(q),g,g_{0}\mid[\alpha,1],[q-1,1],[\beta,1]]$ $q\geq 8$
8i $[\textrm{Sz}(q),g,g_{0}\mid[\alpha,1],[q-1,2],[\beta,1]]$ $q\geq 8$
9 $[{{}^{3}}\textrm{D}_{4}(q),g,g_{0}\mid[q^{4}-q^{2}+1,1]]$
10 $[{{}^{2}}\textrm{G}_{2}(q),g,g_{0}\mid[\frac{q-1}{2},1]]$ $q\geq 27$
11 $[\textrm{M}_{11},g,g_{0}\mid[5,1]]$
12a $[\textrm{M}_{22},g,g_{0}\mid[5,1]]$
12b $[\textrm{M}_{22},g,g_{0}\mid[5,1],[7,1]]$
13 $[\textrm{J}_{1},g,g_{0}\mid[15,1]]$
@@
In Lines 4a and 4b we see the group $\textrm{PSU}_{4}(3)$, and we first look at Line 4a with the list $[\textrm{PSU}_{4}(3),g,g_{0}\mid[5,1]]$. This is a potential branching datum for the group $\textrm{PSU}_{4}(3)$, acting as a group of automorphisms on a compact Riemann surface with genus $g$ and cogenus $g_{0}$ and such that there is exactly one non-regular orbit. The point stabilisers have order $5$ and fix exactly four points, all in one orbit. In Line 4b with the list $[\textrm{PSU}_{4}(3),g,g_{0}\mid[5,1],[7,1]]$, we have the same group, but with two non-regular orbits. On one of them the group acts with fixity 4, as just described, but on the second non-regular orbit we have a fixity three action, with point stabilisers of order $7$. Still, the group acts with fixity $4$ in total, and the difference is that the second possibility exhibits a mixed fixity action.
For the proof of our main theorem, we recall some relevant definitions, and then we refer to Theorem 1.3 of [2] for the specific groups to consider. In the analysis we follow the strategy developed in [19]: First, in Section 3, we analyse series of groups and find generating sets with specific properties. Then we explain the arguments for the proof of Theorem 1.1 in Section 4 and, since some individual groups can be treated with GAP ([6]), we discuss the code that we use in the last section of this article.
Acknowledgements. We remember Kay Magaard with much appreciation and gratitude for initiating the project on groups acting with low fixity and for suggesting questions along the lines of this article.
@@ 2 Preliminaries
If $G$ is a group and $g,h\in G$, then we write conjugation from the right, so $g^{h}:=h^{-1}gh$, and we write commutators as $[g,h]:=g^{-1}h^{-1}gh$.
@@
$\textrm{Alt}_{7}$ $5$ 3 (7)
$\textrm{PSL}_{2}(7)\cong\textrm{PSL}_{3}(2)$ $2$ 2 (3) and 3 (7)
$\textrm{PSL}_{2}(8)$ $2$ 2 (7 or 9)
$\textrm{PSU}_{4}(3)$ $5$ 3 (7)
@@
$\textrm{Sz}(q)$ $q+\sqrt{2q}+1$ or $q-\sqrt{2q}+1$ 2 ($q-1$)
@@ 3 Analysis of series of groups
Referring to (c) we often denote the set as a generating tuple for $G$, namely as $((a_{1},...,a_{g_{0}}),(b_{1},...,b_{g_{0}}),(c_{1,1},...,c_{r,n_{r}}))$, and in the special case of cogenus 00 we write $((~{}),(~{}),(c_{1,1},...,c_{r,n_{r}}))$.
@@
We let $i\in\{1,2\}$ and $j:=3-i$. Since $d_{121}\neq 0\neq d_{212}$ by hypothesis, we find elements $a\in C_{i}$, $b\in C_{j}$ and $g\in G$ in such a way that $a^{g}\cdot b=a$. Then
@@
Therefore the tuple $((a),(g),(b))$ satisfies (a) and (b) for the list $[G,g,1\mid[m_{1},1]]$ if $i=2$, and for the list $[G,g,1\mid[m_{2},1]]$ if $i=1$.
If we look at the list $[G,\bar{g},0\mid[m_{1},2],[m_{2},1]]$ and the case $i=1$, then the tuple $((),(),(a^{-1},a^{g},b))$ satisfies (a) and (b).
Finally, for the list $[G,\bar{g},0\mid[m_{1},1],[m_{2},2]]$ and for the case $i=2$, we see that the tuple $((),(),(b^{-1},(a^{g})^{-1},a))$ satisfies (a) and (b) of Lemma 3.1.
Next we look at the case where $n_{1}+n_{2}=2$. Since $d_{121}\neq 0$, we find $a,c_{1}\in C_{1}$ and $c_{2}\in C_{2}$ such that $c_{1}\cdot c_{2}=a$. Also, we have $d_{212}\neq 0$ and hence there exist $x\in C_{2}$ und $g\in G$ such that $x^{g}\cdot a=x$ (and $x^{-1}\cdot x^{g}=a^{-1}$). We calculate that
@@
showing that the tuple $((x),(g),(c_{1},c_{2}))$ satisfies (a) and (b) from Lemma 3.1 for the list $[G,g,1\mid[m_{1},1],[m_{2},1]]$.
@@
Let $c_{1,1}\in G$ be such that $o(c_{1,1})=\alpha$ and take a subgroup $I\leq G$ of order $\beta$ that contains $c_{1,1}$. Let $t$ denote the central involution in the dihedral group $N:=N_{G}(\langle c_{1,1}\rangle)=N_{G}(I)$ and let $x,y\in G$ be involutions such that $N=\langle x,y\rangle$. We note that $N=C_{G}(t)$ has order $4\cdot\alpha$. Some calculations show that $o(xy)=2\cdot\alpha$ and that $[x,y]=(xy)^{2}$ has order $\alpha$. Hence we may choose $x,y$ such that $[x,y]={c_{1,1}}^{-1}$. In particular $x,y\notin Z(N)$ and we can choose $z\in C_{G}(x)\setminus N$. Then, if we set $a_{1}:=x$ und $b_{1}:=zy$, we see that ${c_{1,1}}^{-1}=[x,y]=[x,zy]=[a_{1},b_{1}]$ and that $((a_{1}),(b_{1}),(c_{1,1}))$ satisfies Lemma 3.1(a) and (b) for the list $l$. We turn to (c) and let $T:=\langle a_{1},b_{1},c_{1,1}\rangle$.
@@
Suppose that $p$ is a prime, $f\in\mathbb{N}$, let $q:=p^{f}\geq 4$, $d:=(q-1,2)$ and $\alpha:=\frac{q+1}{d}$. If $p=2$, then suppose that $f$ is even. Let $H:=\textrm{PSL}_{2}(q)$. Then there are elements $x,y,z\in H$ that generate $H$ and such that $o(x)=2$, $o(z)=\alpha$ and $[x,y]=z$.
@@
We use the subgroup structure of $H$, see for example Theorem 6.5.1 in [7]. Let $z\in H$ be an element of order $\alpha$. Then $N_{H}(\langle z\rangle)$ is dihedral of order $2\alpha$. If we let $C_{1}$ denote a conjugacy class of involutions and $C_{2}:=z^{H}$, then we can calculate (for example with [4]) that the set $M:=\{(a,b)\in C_{1}\times H\mid[a,b]=z^{-1}\}$ has exactly $\alpha\cdot\left|C_{H}(t)\right|$ elements. Now we find $(t,b)\in M$ such that $\langle t,b,z\rangle=H$: If $q\in\{4,5,7,9\}$, then we can check this using GAP. If $q\notin\{4,5,7,9\}$, then $\alpha>5$ and $U:=N_{H}(\langle z\rangle)$ has order at least $12$ and is a maximal subgroup of $H$. Since $[t,b]=1\neq z$, we see that $t$ is not central in $U$, but that it inverts $z$. Now $\left|C_{H}(t)\cap U\right|\in\{2,4\}$. Moreover $\left|C_{H}(t)\right|\geq 8$ because $q\geq 11$, and we choose $g\in C_{H}(t)\setminus U$. Then $(t,g\cdot b)\in M$ and we calculate that $V:=\langle t,g\cdot b,z\rangle$ is in fact all of $H$.
@@
Suppose that $p$ is prime, $f\in\mathbb{N}$ and let $q:=p^{f}\geq 4$, $d:=(q-1,2)$ and $\alpha:=\frac{q+1}{d}$. If $p=2$, then suppose that $f$ is even. Suppose that $G\cong\textrm{PSL}_{2}(q).C_{2}$. Then there are elements $x,y,z\in G$ such that $o(x)=2$, $o(z)=\alpha$, $[x,y]=z$ and $\langle x,y,z\rangle=G$ and such that the following is true:
@@
First suppose that $p\neq 2$. Then $\frac{q+1}{2}=\alpha$ or $\frac{q-1}{2}$ is even and $C_{H}(x)$ is dihedral of order $q+1$ or $q-1$. If $u\in C_{H}(x)$ is an element of order $\alpha$, then $N_{H}(\langle u\rangle)=C_{H}(x)$.Let $U:=\langle g\rangle$ and $\Delta:=U^{G}$. Then all elements of $\Delta$ are subgroups of $H$, because $H\unlhd G$, and then $H$ acts transitively on $\Delta$ by conjugation. If $V\in\Delta$, then its point stabiliser in this action is $N_{G}(V)$ and Frattini gives $G=H\cdot N_{G}(U)$. Moreover
@@
Next suppose that $p=2$. Then $S:=C_{H}(x)$ is elementary abelian of order $q$ and it is a Sylow 2-subgroup of $H$. Moreover $H$ acts transitively on the set of subgroups of order 2 of $S$. Since $H\unlhd G$ gilt, we know that all conjugates of $x$ in $G$ are contained in $H$ and Frattini gives that $G=H\cdot N_{G}(\langle x\rangle)=H\cdot S$. Again we see that $|C_{G}(x):C_{H}(x)|=2$.
For all $p$, we use that $\left|C_{G}(x):C_{H}(x)\right|=2$ and we choose $g\in C_{G}(x)\setminus H$ and set $w:=gy$. Then $1=[x,y]z=[x,w]z$ and we assume for a contradiction that $U:=\langle x,w,z\rangle\neq G$. Since $w\notin H$, this assumption forces $U\neq H$. Moreover $U\neq\langle x,z\rangle$, because $\langle x,z\rangle\leqslant H$ gilt. Let $a\in U\setminus H$. Since $G/H=\{H,Ha\}$, we see that $a\cdot w=agy\in U\cap H$, and we may suppose that $agy\notin\langle x,z\rangle$. Then we choose $a$ such that $H=\langle x,z,agy\rangle\leqslant U$ and it follows that $y\in\langle x,z,agy\rangle=H$ and $y\in U$. Then $g\in U$, contrary to our assumption.
@@
Following the hypotheses in (4.2) and (7.6) of [1], we have that $21\leq l\leq\frac{n}{2}=$ 2 is even, hence $l=2$. We set $C_{l}:=C_{\textrm{SL}(V)}(t)$ and $C:=C_{l}\cap\textrm{Sp}(V)=C_{G}(t)$. Then (4.2) in [1, p. 9] gives that we can choose $t$ to be
@@
We apply Satz II.7.1 in [9] to see that $C$ is a subgroup of a Sylow 2-subgroup $S$ of $\textrm{SL}(V)$. Here $\left|S\right|=q^{6}$ and the matrices of $S$ have arbitrary entries below the main diagonal. We compare this with the entries of matrices from $C$ and the relationships between them, and then it follows that $|C|=\frac{\left|S\right|}{q^{2}}=q^{4}$. We compare this to $\left|G\right|$ (see for example (3.22) in [24, p. 60]) and see that $C\in\textrm{Syl}_{2}(G)$.

### v2

# journal: Journal of Algebra, Computational Section

@@ Abstract.
In previous work, all finite simple groups that act with fixity 4 have been classified. In this article we investigate which ones of these groups act faithfully on a compact Riemann surface of genus at least $2$ with fixity four in total and in such a way that fixity 4 is exhibited on at least one orbit. This is an extended version of the submitted article, including our GAP code.
@@ 1 Introduction
In the following table, $q$ always denotes a prime power. We also remark that, in particular, $g_{0}\geq 0$ if a Hurwitz datum has at least three branchings, and otherwise $g_{0}\geq 1$. But there are some exceptions which we mention in the ”Remark” column.
In Lines 8a–8i we set $\alpha:=q^{2}+\sqrt{2\cdot q}+1$ and $\beta:=q^{2}-\sqrt{2\cdot q}+1$, and $q$ is a power of $2$. The column ”Fixity” indicates the fixity with which $G$ acts on the non-regular orbits, in the order in which they appear in the Hurwitz datum.
@@
Line Hurwitz datum Remark Fixity
1a $\textrm{Alt}_{7},g,g_{0}\mid[5,1]]$ 4
1b $\textrm{Alt}_{7},g,g_{0}\mid[5,1],[7,1]]$ 4, 3
2a $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1]]$ $g_{0}\geq 2$ 4
2b $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,1]]$ 4, 2
2c $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,2]]$ $g_{0}\geq 1$ 4, 2
2d $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[7,1]]$ 4, 3
2e $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,1],[7,1]]$ 4, 3, 2
2f $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,2],[7,1]]$ 4, 3, 2
3a $\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1]]$ $g_{0}\geq 2$ 4
3b $\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,1]]$ 4, 2
3c $\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,2]]$ 4, 2
3d $\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[9,1]]$ 4, 2
3e $\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[9,2]]$ 4, 2
3f $\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,1],[9,1]]$ 4, 2, 2
3g $\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,2],[9,1]]$ 4, 2, 2
3h $\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,1],[9,2]]$ 4, 2, 2
3i $\textrm{PSL}_{2}(8),g,g_{0}\mid[2,1],[7,2],[9,2]]$ 4, 2, 2
4a $\textrm{PSU}_{4}(3),g,g_{0}\mid[5,1]]$ 4
4b $\textrm{PSU}_{4}(3),g,g_{0}\mid[5,1],[7,1]]$ 4, 3
5a $\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q-1}{4},1]]$ $q\geq 9$, $q\equiv 1\mod 4$ 4
5b $\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q+1}{4},1]]$ $q\geq 9$, $q\equiv-1\mod 4$ 4
5c $\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q-1}{4},1],[\frac{q+1}{2},1]]$ $q\geq 9$, $q\equiv 1\mod 4$ 4, 2
5d $\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q-1}{4},1],[\frac{q+1}{2},2]]$ $q\geq 9$, $q\equiv 1\mod 4$ 4, 2
5e $\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q+1}{4},1],[\frac{q-1}{2},1]]$ $q\geq 9$, $q\equiv-1\mod 4$ 4, 2
5f $\textrm{PSL}_{2}(q),g,g_{0}\mid[\frac{q+1}{4},1],[\frac{q-1}{2},2]]$ $q\geq 9$, $q\equiv-1\mod 4$ 4, 2
6 $\textrm{PSp}_{4}(q),g,g_{0}\mid[\frac{q^{2}+1}{(2,q^{2}-1)},1]]$ $q\geq 3$ 4
7 $\textrm{P$\Omega$}^{-}_{8}(q),g,g_{0}\mid[\frac{q^{4}+1}{(2,q^{4}-1)},1]]$ 4
8a $\textrm{Sz}(q),g,g_{0}\mid[\alpha,1]]$ $q\geq 8$ 4
8b $\textrm{Sz}(q),g,g_{0}\mid[\beta,1]]$ $q\geq 8$ 4
8c $\textrm{Sz}(q),g,g_{0}\mid[\alpha,1],[\beta,1]]$ $q\geq 8$ 4, 4
8d $\textrm{Sz}(q),g,g_{0}\mid[\alpha,1],[q-1,1]]$ $q\geq 8$ 4, 2
8e $\textrm{Sz}(q),g,g_{0}\mid[\alpha,1],[q-1,2]]$ $q\geq 8$ 4, 2
8f $\textrm{Sz}(q),g,g_{0}\mid[\beta,1],[q-1,1]]$ $q\geq 8$ 4, 2
8g $\textrm{Sz}(q),g,g_{0}\mid[\beta,1],[q-1,2]]$ $q\geq 8$ 4, 2
8h $\textrm{Sz}(q),g,g_{0}\mid[\alpha,1],[q-1,1],[\beta,1]]$ $q\geq 8$ 4, 2, 4
8i $\textrm{Sz}(q),g,g_{0}\mid[\alpha,1],[q-1,2],[\beta,1]]$ $q\geq 8$ 4, 2, 4
9 ${{}^{3}}\textrm{D}_{4}(q),g,g_{0}\mid[q^{4}-q^{2}+1,1]]$ 4
10 ${{}^{2}}\textrm{G}_{2}(q),g,g_{0}\mid[\frac{q-1}{2},1]]$ $q\geq 27$ 4
11 $\textrm{M}_{11},g,g_{0}\mid[5,1]]$ 4
12a $\textrm{M}_{22},g,g_{0}\mid[5,1]]$ 4
12b $\textrm{M}_{22},g,g_{0}\mid[5,1],[7,1]]$ 4, 3
13 $\textrm{J}_{1},g,g_{0}\mid[15,1]]$ 4
@@
In Lines 4a and 4b we see the group $\textrm{PSU}_{4}(3)$, and we first look at Line 4a with the list $\textrm{PSU}_{4}(3),g,g_{0}\mid[5,1]]$. This is a potential branching datum for the group $\textrm{PSU}_{4}(3)$, acting as a group of automorphisms on a compact Riemann surface with genus $g$ and cogenus $g_{0}$ and such that there is exactly one non-regular orbit. The ”Fixity” column indicates that $G$ acts with fixity $4$ on this orbit, and the Hurwitz datum indicates that the point stabilisers have order $5$ and fix exactly four points, all in this unique non-regular orbit.
In Line 4b with the list $\textrm{PSU}_{4}(3),g,g_{0}\mid[5,1],[7,1]]$, we have the same group, this time with what we call a mixed fixity action: The Hurwitz datum shows us two non-regular orbits, and point stabilisers of order $5$ and order $7$, respectively. Together with the ”Fixity” column, we see that the first non-regular orbit belongs to a fixity $4$ action with point stabilisers of order $5$, and the second non-regular orbit belongs to a fixity $3$ action with point stabilisers of order $7$.
For the proof of our main theorem, we recall some relevant definitions, and then we refer to Theorem 1.3 of [2] for the specific groups to consider. In the analysis we follow the strategy developed in [20]: First, in Section 3, we analyse series of groups and find generating sets with specific properties. Then we explain the arguments for the proof of Theorem 1.1 in Section 4. Some individual groups can be treated with GAP ([6]), and we show and comment on the code at the end of this article.
Acknowledgements. We remember Kay Magaard with much appreciation and gratitude for initiating the project on groups acting with low fixity and for suggesting questions along the lines of this article. We are also grateful to the referee because one of their comments gave us the idea to add information to Table 1.
@@ 2 Preliminaries
If $G$ is a group and $g,h\in G$, then we write conjugation from the right, so $g^{h}:=h^{-1}gh$, and we write commutators as $g,h]:=g^{-1}h^{-1}gh$.
@@
$\textrm{Alt}_{7}$ 55 3 (7)
$\textrm{PSL}_{2}(7)\cong\textrm{PSL}_{3}(2)$ 22 2 (3) and 3 (7)
$\textrm{PSL}_{2}(8)$ 22 2 (7 or 9)
$\textrm{PSU}_{4}(3)$ 55 3 (7)
@@
$\textrm{Sz}(q)$ $q+\sqrt{2q}+1$ or $q-\sqrt{2q}+1$ 2 (q1q-1)
@@ 3 Analysis of series of groups
Referring to (c) we often denote the set as a generating tuple for $G$, namely as $((a_{1},...,a_{g_{0}}),(b_{1},...,b_{g_{0}}),(c_{1,1},...,c_{r,n_{r}}))$, and in the special case of cogenus 0 we write $((~),(~),(c_{1,1},...,c_{r,n_{r}}))$.
@@
We let $i\in\{1,2\}$ and j:=3ij:=3-i. Since $d_{121}\neq 0\neq d_{212}$ by hypothesis, we find elements $a\in C_{i}$, $b\in C_{j}$ and $g\in G$ in such a way that $a^{g}\cdot b=a$. Then
@@
Therefore the tuple $((a),(g),(b))$ satisfies (a) and (b) for the list $G,g,1\mid[m_{1},1]]$ if $i=2$, and for the list $G,g,1\mid[m_{2},1]]$ if $i=1$.
If we look at the list $G,\bar{g},0\mid[m_{1},2],[m_{2},1]]$ and the case $i=1$, then the tuple $((),(),(a^{-1},a^{g},b))$ satisfies (a) and (b).
Finally, for the list $G,\bar{g},0\mid[m_{1},1],[m_{2},2]]$ and for the case $i=2$, we see that the tuple $((),(),(b^{-1},(a^{g})^{-1},a))$ satisfies (a) and (b) of Lemma 3.1.
Next we look at the case where $n_{1}+n_{2}=2$. Since $d_{121}\neq 0$, we find $a,c_{1}\in C_{1}$ and $c_{2}\in C_{2}$ such that $c_{1}\cdot c_{2}=a$. Also, we have $d_{212}\neq 0$ and hence there exist $x\in C_{2}$ and $g\in G$ such that $x^{g}\cdot a=x$ (and $x^{-1}\cdot x^{g}=a^{-1}$). We calculate that
@@
showing that the tuple $((x),(g),(c_{1},c_{2}))$ satisfies (a) and (b) from Lemma 3.1 for the list $G,g,1\mid[m_{1},1],[m_{2},1]]$.
@@
Let $c_{1,1}\in G$ be such that $o(c_{1,1})=\alpha$ and take a subgroup $I\leq G$ of order $\beta$ that contains $c_{1,1}$. Let $t$ denote the central involution in the dihedral group $N:=N_{G}(\langle c_{1,1}\rangle)=N_{G}(I)$ and let $x,y\in G$ be involutions such that $N=\langle x,y\rangle$. We note that $N=C_{G}(t)$ has order $4\cdot\alpha$. Some calculations show that $o(xy)=2\cdot\alpha$ and that $x,y]=(xy)^{2}$ has order $\alpha$. Hence we may choose $x,y$ such that $x,y]={c_{1,1}}^{-1}$. In particular $x,y\notin Z(N)$ and we can choose $z\in C_{G}(x)\setminus N$. Then, if we set $a_{1}:=x$ and $b_{1}:=zy$, we see that ${c_{1,1}}^{-1}=[x,y]=[x,zy]=[a_{1},b_{1}]$ and that $((a_{1}),(b_{1}),(c_{1,1}))$ satisfies Lemma 3.1(a) and (b) for the list $l$. We turn to (c) and let $T:=\langle a_{1},b_{1},c_{1,1}\rangle$.
@@
Suppose that $p$ is a prime, $f\in\mathbb{N}$, let $q:=p^{f}\geq 4$, d:=(q1,2)d:=(q-1,2) and $\alpha:=\frac{q+1}{d}$. If $p=2$, then suppose that $f$ is even. Let $H:=\textrm{PSL}_{2}(q)$. Then there are elements $x,y,z\in H$ that generate $H$ and such that $o(x)=2$, $o(z)=\alpha$ and $[x,y]=z$.
@@
We use the subgroup structure of $H$, see for example Theorem 6.5.1 in [7]. Let $z\in H$ be an element of order $\alpha$. Then $N_{H}(\langle z\rangle)$ is dihedral of order $2\alpha$. If we let $C_{1}$ denote a conjugacy class of involutions and $C_{2}:=z^{H}$, then we can calculate (for example with [4]) that the set $M:=\{(a,b)\in C_{1}\times H\mid[a,b]=z^{-1}\}$ has exactly $\alpha\cdot\left|C_{H}(t)\right|$ elements. Now we find $(t,b)\in M$ such that $\langle t,b,z\rangle=H$: If $q\in\{4,5,7,9\}$, then we can check this using GAP ([6]). If $q\notin\{4,5,7,9\}$, then $\alpha>5$ and $U:=N_{H}(\langle z\rangle)$ has order at least $12$ and is a maximal subgroup of $H$. Since $t,b]=1\neq z$, we see that $t$ is not central in $U$, but that it inverts $z$. Now $\left|C_{H}(t)\cap U\right|\in\{2,4\}$. Moreover $\left|C_{H}(t)\right|\geq 8$ because $q\geq 11$, and we choose $g\in C_{H}(t)\setminus U$. Then $(t,g\cdot b)\in M$ and we calculate that $V:=\langle t,g\cdot b,z\rangle$ is in fact all of $H$.
@@
Suppose that $p$ is prime, $f\in\mathbb{N}$ and let $q:=p^{f}\geq 4$, d:=(q1,2)d:=(q-1,2) and $\alpha:=\frac{q+1}{d}$. If $p=2$, then suppose that $f$ is even. Suppose that $G\cong\textrm{PSL}_{2}(q).C_{2}$. Then there are elements $x,y,z\in G$ such that $o(x)=2$, $o(z)=\alpha$, $[x,y]=z$ and $\langle x,y,z\rangle=G$ and such that the following is true:
@@
First suppose that $p\neq 2$. Then $\frac{q+1}{2}=\alpha$ or $\frac{q-1}{2}$ is even and $C_{H}(x)$ is dihedral of order $q+1$ or q1q-1. If $u\in C_{H}(x)$ is an element of order $\alpha$, then $N_{H}(\langle u\rangle)=C_{H}(x)$.Let $U:=\langle g\rangle$ and $\Delta:=U^{G}$. Then all elements of $\Delta$ are subgroups of $H$, because $H\unlhd G$, and then $H$ acts transitively on $\Delta$ by conjugation. If $V\in\Delta$, then its point stabiliser in this action is $N_{G}(V)$ and Frattini gives $G=H\cdot N_{G}(U)$. Moreover
@@
Next suppose that $p=2$. Then $S:=C_{H}(x)$ is elementary abelian of order $q$ and it is a Sylow 2-subgroup of $H$. Moreover $H$ acts transitively on the set of subgroups of order 2 of $S$. Since $H\unlhd G$, we know that all conjugates of $x$ in $G$ are contained in $H$ and Frattini gives that $G=H\cdot N_{G}(\langle x\rangle)=H\cdot S$. Again we see that $|C_{G}(x):C_{H}(x)|=2$.
For all $p$, we use that $\left|C_{G}(x):C_{H}(x)\right|=2$ and we choose $g\in C_{G}(x)\setminus H$ and set $w:=gy$. Then $1=[x,y]z=[x,w]z$ and we assume for a contradiction that $U:=\langle x,w,z\rangle\neq G$. Since $w\notin H$, this assumption forces $U\neq H$. Moreover $U\neq\langle x,z\rangle$, because $\langle x,z\rangle\leqslant H$. Let $a\in U\setminus H$. Since $G/H=\{H,Ha\}$, we see that $a\cdot w=agy\in U\cap H$, and we may suppose that $agy\notin\langle x,z\rangle$. Then we choose $a$ such that $H=\langle x,z,agy\rangle\leqslant U$ and it follows that $y\in\langle x,z,agy\rangle=H$ and $y\in U$. Then $g\in U$, contrary to our assumption.
@@
Following the hypotheses in (4.2) and (7.6) of [1], we have that $1\leq l\leq\frac{n}{2}=2$ is even, hence $l=2$. We set $C_{l}:=C_{\textrm{SL}(V)}(t)$ and $C:=C_{l}\cap\textrm{Sp}(V)=C_{G}(t)$. Then (4.2) in [1, p. 9] gives that we can choose $t$ to be
@@
We apply Satz II.7.1 in [9] to see that $C$ is a subgroup of a Sylow 2-subgroup $S$ of $\textrm{SL}(V)$. Here $\left|S\right|=q^{6}$ and the matrices of $S$ have arbitrary entries below the main diagonal. We compare this with the entries of matrices from $C$ and the relationships between them, and then it follows that $|C|=\frac{\left|S\right|}{q^{2}}=q^{4}$. We compare this to $\left|G\right|$ (see for example (3.22) in [25, p. 60]) and see that $C\in\textrm{Syl}_{2}(G)$.